# Finite density QCD equation of state: critical point and lattice-based $T^{\prime}$-expansion

### v2

@@ II.1 Taylor Expansion
$\frac{P(T,\mu_{B})}{T^{4}}=\sum_{n=0}{\frac{1}{2n!}\chi_{2n}(T,\mu_{B}=0)\left(\frac{\mu_{B}}{T}\right)^{2n}}\,\,,$ (1)
@@
$\chi_{n}(T)=\left(\frac{\partial^{n}}{\partial(\mu_{B}/T)^{n}}\frac{P(T,\mu_{B})}{T^{4}}\right)_{\mu_{B}=0}\,\,.$
@@
$=$ $\sum_{n=1}^{\infty}\frac{1}{(2n-1)!}\chi_{2n}(T)\left(\frac{\mu_{B}}{T}\right)^{2n-1}.$
@@
$n_{B}(T,\mu_{B})$ $=T^{3}\sum_{n=1}^{2}\frac{1}{(2n-1)!}\chi_{2n}^{\rm non-Ising}(T)\left(\frac{\mu_{B}}{T}\right)^{2n-1}$
@@
where $n_{B}^{\rm Ising}$ is the contribution to the baryon density with the singular behavior appropriate for the 3D Ising critical point, and the coefficients $\chi_{n}^{\rm non-Ising}(T)$ satisfy
$\chi_{n}^{\rm lat}(T)=\chi_{n}^{\rm non-Ising}(T)+\frac{T_{C}^{4}}{T^{4}}\chi_{n}^{\rm Ising}(T)$
for $n=0,~{}2,~{}4$, where $\chi_{n}^{\rm lat}(T)$ are the input from lattice QCD and $\chi_{n}^{\rm Ising}(T)$ represent the critical contribution to the expansion coefficients. Although this construction works well, it was observed that at large values of $\mu_{B}$, wiggles appear in the thermodynamic observables, particularly the baryon density and speed of sound, for some parameter choices. This is due to the truncation in the Taylor expansion of the non-Ising contribution to the observables, which limits the current applicability of this equation of state.
@@ II.2 $T^{\prime}$-Expansion Scheme
where the Taylor expansion coefficients $\kappa_{2}^{BB}$, etc. are almost constant as functions of $T$ in the transition region, while the rapid changes in EoS associated with the crossover are mostly captured by the function $\chi_{2}(T^{\prime},0)$ (see Fig.2 below.).
The “$T^{\prime}$-expansion” scheme is essentially a re-shuffling of the Taylor expansion in Eq.(2), and the coefficients $\kappa_{n}^{BB}(T)$ can be expressed in terms of the susceptibilities $\chi_{2n}(T)$:
@@
These coefficients were obtained in high-statistics lattice QCD simulations [13]. As expected, compared to the sharply rising $\chi_{2}^{B}(T)$, $\kappa_{2}^{BB}$ shows a very mild temperature dependence around the transition region, which makes the $T^{\prime}$-expansion scheme more favorable than the Taylor expansion since it does not introduce the wiggly behavior in the EoS at large $\mu_{B}$. Moreover, the fact that $\kappa_{4}$ is shown in [13] to be consistent with zero hints at a faster convergence compared to the Taylor series.
@@ II.3 Lattice data
For convenience we introduce an auxiliary variable $x=\frac{T}{200\,\text{MeV}}$. For $\chi_{2,\text{lat}}^{B}(T)$, we employ four free parameters $d_{i}$, such that the crossover occurs at $x\approx d_{1}$, and its width $\Delta x\sim d_{1}/d_{2}$ is controlled by $d_{2}\gg 1$, while $d_{3}\left(1-\frac{d_{4}^{2}}{x^{2}}\right)$ provides large-$x$ asymptotics:
@@
where, $m_{p}$ denotes the proton mass (in units of 200 MeV), and the first term, typically very small, yields the correct low-temperature asymptotics for $\chi_{2}^{B}$ in QCD, representing the nonrelativistic contribution of nucleons/antinucleons. Best-fit coefficients for $\chi_{2,\text{lat}}^{B}(T)$ are listed in Table 1, and the resulting parametrization is shown in the top panel of Fig. 2.
@@
For $\kappa_{2}^{BB}$, we employ a rational fit, enforcing the expected small-$x$ linear behaviour and the large-$x$ $\kappa_{2}^{BB}\sim A\,x^{2}$ dependence, with $A=\chi_{4,\rm SB}^{B}/(12d_{3}d_{4}^{2})$, where $\chi_{4,\rm SB}^{B}$ is the fourth baryon susceptibility in the infinite-temperature limit.
$\kappa_{2}^{BB}(T)=\frac{a_{1}x+a_{2}x^{2}+a_{3}x^{3}+a_{4}x^{4}+a_{5}x^{5}}{b_{0}+b_{1}x+b_{2}x^{2}+a_{5}/A\,x^{3}}$ (8)
and again $x=\frac{T}{200\,\rm MeV}$. Best-fit parameters for $\kappa_{2}^{BB}(T)$ are listed in Table 2, and the resulting parametrization is shown in the bottom panel of Fig. 2.
$a_{1}$ $a_{2}$ $a_{3}$ $a_{4}$ $a_{5}$ $b_{0}$ $b_{1}$ $b_{2}$ $b_{3}$
0.48 -0.64 0.52 -0.78 0.53 11.9 -9.62 -7.51 10.8
Table 2: Coefficients of the parameterized form for $\kappa^{B}_{2}(T)$ from Eq (8) for $25~{}\text{MeV}\leq T\leq 800~{}\text{MeV}$ given as a ratio.
@@ III.1 Scaling: 3D Ising Model
In this work, we employ the same form of the scaling equation of state as used in the BEST collaboration equation of state. The parameterization of magnetization, denoted by $M$, reduced temperature ($r$) and external magnetic field ($h$) in terms of additional scaling variables $R$ and $\theta$, is given as follows [51, 75, 76, 77, 78, 79, 80]:
@@
The given parameterization involves an odd function $\tilde{h}(\theta)=\theta(1+a\theta^{2}+b\theta^{4})$, where $a=-0.76201$ and $b=0.00804$. The critical exponents for the 3D Ising model are $\beta=0.326$ and $\delta=4.80$. It’s worth noting that $R$ is required to be non-negative ($R\geq 0$), and $|\theta|$ should be less than or equal to the first non-trivial zero of $\tilde{h}(\theta)$, denoted as $\theta_{0}\simeq 1.154$. To fix the values of the normalization constants $M_{0}$ and $h_{0}$, two conditions $M(r=-1,h=0^{+})=1$ and $M(r=0,h=1)\propto\text{sgn}(h)|h|^{1/\delta}$ are used. These conditions result in $M_{0}=0.605$ and $h_{0}=0.394$. It is important to note that this parametric representation gives a non-globally invertible mapping from $(R,\theta)\mapsto(r,h)$. The critical point is located at $(r=0,h=0)$, and when $r<0$, there is a smooth transition (crossover), while $r>0$ corresponds to a first-order phase transition.
@@ III.2 Mapping 3D Ising coordinates to QCD coordinates
The parameters $(w,\rho)$ act as scaling factors for the variables $r$ and $h$, where $w$ determines the size of the critical region, and $\rho$ modifies its shape. The difference $\alpha_{12}$ between $\alpha_{2}$ and $\alpha_{1}$ also controls the strength of the discontinuity. Equations (III.2) can be inverted to give:
@@ III.3 Transition Line
$\chi_{1}^{\rm crit}=\frac{n_{B}^{\rm crit}(T,\mu_{B})}{T^{3}}=\frac{\partial(P^{\rm crit}(T,\mu_{B})/T^{4})}{\partial(\mu_{B}/T)}\Big{|}_{T}\,\,.$ (20)
@@
In this paper, we illustrate two choices of critical baryon chemical potential. The first one, used mainly for comparison with the BEST collaboration EoS, is $\mu_{BC}=350~{}\text{MeV}$, giving $T_{C}=140~{}\text{MeV}$ and $\alpha_{1}=6.7^{\circ}$. For the first choice of parameters, we show in Fig. 4 contours of equal normalized critical pressure, in the $r-h$, $T^{\prime}-\mu_{B}^{2}$ and $T-\mu_{B}$ planes. The first order transition line is shown as a red solid line, and the critical point corresponds to a black dot. We show positive and negative values of $\mu_{B}$, corresponding to positive and negative baryon chemical potentials, to illustrate the symmetry of QCD under baryon-and-antibaryon exchange. This symmetry arises naturally from the selection of a quadratic mapping of the chemical potential in Eq. (13). For the same choice of parameters, in Fig. 5 we show the critical baryon density, which develops a discontinuity for $\mu_{B}>\mu_{BC}$, as required for a first order transition. With the second choice we place the critical point in a region which goes beyond the limits of the BEST collaboration EoS: we choose $\mu_{BC}=500~{}\text{MeV}$, corresponding to $T_{C}=116~{}\text{MeV}$ and $\alpha_{1}=11.2^{\circ}$.
@@ VI.3 Stability and Causality
must fall within the range $0\leq c_{s}^{2}\leq 1$. The behavior of the speed of sound as a function of $T$ and $\mu_{B}$ can be seen in Fig. 11. It exhibits a dip at the critical point, where it vanishes. We show in Fig. 12 a landscape of acceptable (blue triangles) and pathological (red squares) choices for the parameters $w$ and $\rho$, for a critical point located at $\mu_{BC}=500~{}\text{MeV}$, which corresponds to $T_{C}=117~{}\text{MeV}$ and $\alpha_{1}=11^{\circ}$. Additionally, we have $\alpha_{12}=\alpha_{1}$, meaning $\alpha_{2}=\alpha_{1}-\alpha_{12}$, while $w$ and $\rho$ are varied in the range $w=2.5-22.5$, $\rho=0.1-1.3$. Similar plots, comparing our parameter landscapes to the ones from the BEST collaboration EoS, are discussed in Appendix C.
@@ Acknowledgements.
This material is based upon work supported by the National Science Foundation under grants No. PHY-2208724, PHY-1654219 and PHY-2116686, and within the framework of the MUSES collaboration, under grant number No. OAC-2103680. This material is also based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under Awards Number DE-SC0022023, DE-FG02-05ER41367 and DE-FG0201ER41195. O.S. and E.B. acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the grant CRC-TR 211 ’Strong-interaction matter under extreme conditions’ - Project number 315477589 - TRR 211. This work is supported by the European Union’s Horizon 2020 research and innovation program under grant agreement No 824093 (STRONG-2020).
@@ Appendix A Taylor expansion of the critical contribution
Here we present the formula and derivation for $\text{Taylor}[T^{\prime}_{\rm crit}(T,\mu_{B})]$. Since in our approach $T^{\prime}_{\rm lat}(T,\mu_{B})$ is truncated up to $\kappa_{2}(T)$, we need the $\text{Taylor}[T^{\prime}_{crit}(T,\mu_{B}),n=2]=a_{0}(T)+a_{2}(T)\left(\frac{\mu_{B}}{T}\right)^{2}$, such that we match that same order by construction, while the higher order contributions come from the critical part. The coefficients $a_{0}$ and $a_{2}$ are then given by;
$a_{0}(T)$ $=\left(\frac{\partial\chi_{2}}{\partial T}\bigg{|}_{T_{0}}\right)\left(\frac{\partial n_{B}^{\rm crit}(T,\mu_{B})}{\partial(\mu_{B}/T)}\right)\bigg{|}_{\mu_{B}/T=0}$ (32)
$a_{2}(T)$ $=\left(\frac{\partial\chi_{2}}{\partial T}\bigg{|}_{T_{0}}\right)\left(\frac{1}{3!}\frac{\partial^{3}n_{B}^{\rm crit}(T,\mu_{B})}{\partial(\mu_{B}/T)^{3}}\right)\bigg{|}_{\mu_{B}/T=0}$ (33)
@@ Appendix B Computing thermodynamics
$\frac{\mu_{B}}{T}\frac{\partial\chi_{2,\rm lat}(T)}{\partial T}\frac{\partial T^{\prime}}{\partial T}T^{3}$ (35)
@@ Appendix C Comparison with BEST EoS
and linearizing Eq. (13) around the critical point and using $T^{\prime}_{\text{lat}}=T\left(1+\kappa_{2}(T)\left(\frac{\mu_{B}}{T}\right)^{2}\right)$,
$\frac{1}{T^{\prime}_{,T}}\frac{\Delta T^{\prime}}{\Delta\mu_{B}}=\frac{\Delta T}{\Delta\mu_{B}}+\frac{2\kappa_{2}(T)\mu_{B}}{T^{\prime}_{,T}T}$ (40)
@@
$\tan\alpha_{1}=\frac{2\kappa_{2}(T_{C})\mu_{BC}}{T^{\prime}_{,T}T_{C}}$ (41)

### v3

@@ II.1 Taylor Expansion
$\frac{P(T,\mu_{B})}{T^{4}}=\sum_{n=0}{\frac{1}{2n!}\chi_{2n}^{B}(T,\mu_{B}=0)\left(\frac{\mu_{B}}{T}\right)^{2n}}\,\,,$ (1)
@@
$\chi_{n}^{B}(T)=\left(\frac{\partial^{n}}{\partial(\mu_{B}/T)^{n}}\frac{P(T,\mu_{B})}{T^{4}}\right)_{\mu_{B}=0}\,\,.$
@@
$=$ $\sum_{n=1}^{\infty}\frac{1}{(2n-1)!}\chi_{2n}^{B}(T)\left(\frac{\mu_{B}}{T}\right)^{2n-1}.$
@@
$n_{B}(T,\mu_{B})$ $=T^{3}\sum_{n=1}^{2}\frac{1}{(2n-1)!}\chi_{2n}^{B~{}\rm non-Ising}(T)\left(\frac{\mu_{B}}{T}\right)^{2n-1}$
@@
where $n_{B}^{\rm Ising}$ is the contribution to the baryon density with the singular behavior appropriate for the 3D Ising critical point, and the coefficients $\chi_{n}^{B~{}\rm non-Ising}(T)$ satisfy
$\chi_{n}^{B~{}\rm lat}(T)=\chi_{n}^{B~{}\rm non-Ising}(T)+\frac{T_{C}^{4}}{T^{4}}\chi_{n}^{B~{}\rm Ising}(T)$
for $n=0,~{}2,~{}4$, where $\chi_{n}^{B~{}\rm lat}(T)$ are the input from lattice QCD and $\chi_{n}^{B~{}\rm Ising}(T)$ represent the critical contribution to the expansion coefficients. Although this construction works well, it was observed that at large values of $\mu_{B}$, wiggles appear in the thermodynamic observables, particularly the baryon density and speed of sound, for some parameter choices. This is due to the truncation in the Taylor expansion of the non-Ising contribution to the observables, which limits the current applicability of this equation of state.
@@ II.2 $T^{\prime}$-Expansion Scheme
where the Taylor expansion coefficients $\kappa_{2}^{BB}$, etc. are almost constant as functions of $T$ in the transition region, while the rapid changes in EoS associated with the crossover are mostly captured by the function $\chi_{2}^{B}(T^{\prime},0)$ (see Fig.2 below.).
The “$T^{\prime}$-expansion” scheme is essentially a re-shuffling of the Taylor expansion in Eq.(2), and the coefficients $\kappa_{n}^{BB}(T)$ can be expressed in terms of the susceptibilities $\chi_{2n}^{B}(T)$:
@@
These coefficients were obtained in high-statistics lattice QCD simulations [13]. As expected, compared to the sharply rising $\chi_{2}^{B}(T)$, $\kappa_{2}^{BB}$ shows a very mild temperature dependence around the transition region, which makes the $T^{\prime}$-expansion scheme more favorable than the Taylor expansion since it does not introduce the wiggly behavior in the EoS at large $\mu_{B}$. Moreover, the fact that $\kappa_{4}^{BB}$ is shown in [13] to be consistent with zero hints at a faster convergence compared to the Taylor series.
@@ II.3 Lattice data
For convenience we introduce an auxiliary variable $x={T}/({200\,\text{MeV}})$. For $\chi_{2,\text{lat}}^{B}(T)$, we employ four free parameters $d_{i}$, such that the crossover occurs at $x\approx d_{1}$, and its width $\Delta x\sim d_{1}/d_{2}$ is controlled by $d_{2}\gg 1$, while $d_{3}\left(1-\frac{d_{4}^{2}}{x^{2}}\right)$ provides large-$x$ asymptotics:
@@
where, $m_{p}\approx 4.7$ denotes the proton mass (in units of 200 MeV). The first term, typically very small, yields the correct low-temperature asymptotics for $\chi_{2}^{B}$ in QCD, representing the nonrelativistic contribution of nucleons/antinucleons. Best-fit coefficients for $\chi_{2,\text{lat}}^{B}(T)$ are listed in Table 1, and the resulting parametrization is shown in the top panel of Fig. 2.
@@
Because $\kappa_{2}^{BB}$, unlike $\chi_{2}^{B}$, varies slowly in the temperature region of our interest, we can use a rational fit. We enforce the expected small-temperature linear behaviour which follows from the dominant exponential behavior $\chi^{B}_{4},\chi^{B}_{2}\sim e^{-m_{p}/x}$ and Eq. 6: $\kappa_{2}^{BB}/x\to A_{1}=1/(6m_{p})\approx 0.035$ Similarly, the leading large-temperature behavior follows from Eq.(7), i.e., $\chi_{2}=d_{3}-d_{3}d_{4}^{2}/x^{2}+\dots$ and $\chi_{4}=2/({9\pi^{2}})+\dots$. Eq. 6 then gives $\kappa_{2}^{BB}/x^{2}\to A_{2}=1/(54d_{3}d_{4}^{2})\approx 1.47$. We then use the following fitting function:
$\kappa_{2}^{BB}(T)=\frac{A_{1}b_{0}x+a_{2}x^{2}+a_{3}x^{3}+A_{2}x^{4}}{b_{0}+b_{1}x+x^{2}}$ (8)
where, again, $x={T}/({200\,\rm MeV})$. Best-fit parameters for $\kappa_{2}^{BB}(T)$ are listed in Table 2, and the resulting parametrization is shown in the bottom panel of Fig. 2.
$a_{2}$ $a_{3}$ $b_{0}$ $b_{1}$
0.652 -2.60 21.4 -9.81
Table 2: Coefficients of the rational parameterization for $\kappa^{BB}_{2}(T)$ in Eq (8) for $25~{}\text{MeV}\leq T\leq 800~{}\text{MeV}$.
@@ III.1 Scaling: 3D Ising Model
In this work, we employ the same form of the scaling equation of state as used in the BEST collaboration equation of state. The parameterization of magnetization, denoted by $M$, reduced temperature ($r$) and external magnetic field ($h$) in terms of additional scaling standard variables $R$ and $\theta$, is given as follows [51, 75, 76, 77, 78, 79, 80]:
@@
The scale invariant “angular” variable $\theta$ describes the position of a point on $r,h$ plane relative to the $h=0$ ($\theta=0$) and $r=0$ ($\theta=1$) axes, allowing a non-singular description of both regimes for $R\neq 0$. The “radial” variable $R$ measures the distance from the critical point, $R=0$. The parameterization involves an odd function $\tilde{h}(\theta)=\theta(1+a\theta^{2}+b\theta^{4})$, where $a=-0.76201$ and $b=0.00804$. The critical exponents for the 3D Ising model are $\beta=0.326$ and $\delta=4.80$. While $R$ is non-negative ($R\geq 0$), $|\theta|$ should not exceed the first non-trivial zero of $\tilde{h}(\theta)$, denoted as $\theta_{0}\simeq 1.154$ and corresponding to $r<0$, $h=0$ axis. To fix the values of the normalization constants $M_{0}$ and $h_{0}$, two conditions $M(r=-1,h=0^{+})=1$ and $M(r=0,h=1)=1$ are used. These conditions result in $M_{0}=0.605$ and $h_{0}=0.364$. It is important to note that this parametric representation gives a non-globally invertible mapping from $(R,\theta)\mapsto(r,h)$. The critical point is located at $(r=0,h=0)$, and when $r<0$, there is a smooth transition (crossover), while $r>0$ corresponds to a first-order phase transition.
@@ III.2 Mapping 3D Ising coordinates to QCD coordinates
The parameters $(w,\rho)$ act as scaling factors for the variables $r$ and $h$, where $w$ determines the size of the critical region, and $\rho$ modifies its shape. The difference $\alpha_{12}$ between $\alpha_{2}$ and $\alpha_{1}$ also controls the strength of the discontinuity. Equations (14) can be inverted to give:
@@ III.3 Transition Line
$\chi_{1}^{B~{}\rm crit}=\frac{n_{B}^{\rm crit}(T,\mu_{B})}{T^{3}}=\frac{\partial(P^{\rm crit}(T,\mu_{B})/T^{4})}{\partial(\mu_{B}/T)}\Big{|}_{T}\,\,.$ (20)
@@
In this paper, we illustrate two choices of critical baryon chemical potential. The first one, used mainly for comparison with the BEST collaboration EoS, is $\mu_{BC}=350~{}\text{MeV}$, giving $T_{C}=140~{}\text{MeV}$ and $\alpha_{1}=6.7^{\circ}$. For the first choice of parameters, we show in Fig. 4 contours of equal normalized critical pressure, in the $r-h$, $T^{\prime}-\mu_{B}^{2}$ and $T-\mu_{B}$ planes. The first-order transition line is shown as a red solid line, and the critical point corresponds to a black dot. We show positive and negative values of $\mu_{B}$, corresponding to positive and negative baryon chemical potentials, to illustrate the symmetry of QCD under baryon-and-antibaryon exchange. This symmetry arises naturally from the selection of a quadratic mapping of the chemical potential in Eq. (13). For the same choice of parameters, in Fig. 5 we show the critical baryon density, which develops a discontinuity for $\mu_{B}>\mu_{BC}$, as required for a first order transition. With the second choice we place the critical point in a region that goes beyond the limits of the BEST collaboration EoS: we choose $\mu_{BC}=500~{}\text{MeV}$, corresponding to $T_{C}=116~{}\text{MeV}$ and $\alpha_{1}=11.2^{\circ}$.
@@ VI.3 Stability and Causality
must fall within the range $0\leq c_{s}^{2}\leq 1$. The behavior of the speed of sound as a function of $T$ and $\mu_{B}$ can be seen in Fig. 11. It exhibits a dip at the critical point, where it vanishes. We show in Fig. 12 a landscape of acceptable (blue dots) and pathological (red squares) choices for the parameters $w$ and $\rho$, for a critical point located at $\mu_{BC}=500~{}\text{MeV}$, which corresponds to $T_{C}=117~{}\text{MeV}$ and $\alpha_{1}=11^{\circ}$. Additionally, we have $\alpha_{12}=\alpha_{1}$, meaning $\alpha_{2}=\alpha_{1}-\alpha_{12}$, while $w$ and $\rho$ are varied in the range $w=2.5-22.5$, $\rho=0.1-1.3$. Similar plots, comparing our parameter landscapes to the ones from the BEST collaboration EoS, are discussed in Appendix C.
@@ Acknowledgements.
This material is based upon work supported by the National Science Foundation under grants No. PHY-2208724, PHY-1654219 and PHY-2116686, within the framework of the MUSES collaboration, under grant number No. OAC-2103680 and by the National Aeronautics and Space Agency (NASA) under Award Number 80NSSC24K0767. This material is also based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under Awards Number DE-SC0022023, DE-FG02-05ER41367 and DE-FG0201ER41195. O.S. and E.B. acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the grant CRC-TR 211 ’Strong-interaction matter under extreme conditions’ - Project number 315477589 - TRR 211. This work is supported by the European Union’s Horizon 2020 research and innovation program under grant agreement No 824093 (STRONG-2020).
@@ Appendix A Taylor expansion of the critical contribution
Here we present the formula and derivation for $\text{Taylor}[T^{\prime}_{\rm crit}(T,\mu_{B})]$. Since in our approach $T^{\prime}_{\rm lat}(T,\mu_{B})$ is truncated up to $\kappa_{2}^{BB}(T)$, we need the $\text{Taylor}[T^{\prime}_{crit}(T,\mu_{B}),n=2]=a_{0}(T)+a_{2}(T)\left(\frac{\mu_{B}}{T}\right)^{2}$, such that we match that same order by construction, while the higher order contributions come from the critical part. The coefficients $a_{0}$ and $a_{2}$ are then given by;
$a_{0}(T)$ $=\left(\frac{\partial\chi_{2}^{B}}{\partial T}\bigg{|}_{T_{0}}\right)\left(\frac{\partial n_{B}^{\rm crit}(T,\mu_{B})}{\partial(\mu_{B}/T)}\right)\bigg{|}_{\mu_{B}/T=0}$ (32)
$a_{2}(T)$ $=\left(\frac{\partial\chi_{2}^{B}}{\partial T}\bigg{|}_{T_{0}}\right)\left(\frac{1}{3!}\frac{\partial^{3}n_{B}^{\rm crit}(T,\mu_{B})}{\partial(\mu_{B}/T)^{3}}\right)\bigg{|}_{\mu_{B}/T=0}$ (33)
@@ Appendix B Computing thermodynamics
$\frac{\mu_{B}}{T}\frac{\partial\chi_{2,\rm lat}^{B}(T)}{\partial T}\frac{\partial T^{\prime}}{\partial T}T^{3}$ (35)
@@ Appendix C Comparison with BEST EoS
and linearizing Eq. (13) around the critical point and using $T^{\prime}_{\text{lat}}=T\left(1+\kappa_{2}^{BB}(T)\left(\frac{\mu_{B}}{T}\right)^{2}\right)$,
$\frac{1}{T^{\prime}_{,T}}\frac{\Delta T^{\prime}}{\Delta\mu_{B}}=\frac{\Delta T}{\Delta\mu_{B}}+\frac{2\kappa_{2}^{BB}(T)\mu_{B}}{T^{\prime}_{,T}T}$ (40)
@@
$\tan\alpha_{1}=\frac{2\kappa_{2}^{BB}(T_{C})\mu_{BC}}{T^{\prime}_{,T}T_{C}}$ (41)